Find the t_a/2 value for a 95% confidence interval when the sample size is 22.

1
Expert's answer
2022-02-21T14:13:47-0500

For a confidence level of 95%, we know that "\alpha = 1 - 0.95 = 0.05" and a sample size of "n = 22," we get "df = 22-1 =21" degrees of freedom, and using a t-distribution table we find that 

"t_{\alpha/2, n-1}=t_{0.025, 21}= 2.079614"